OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x

OpenStudy (anonymous):

sin3x = sin(2x+x) = sin(2x)cosx + cos(2x) sinx = .... More hints: sin2x = 2sinxcosx cos2x = 1-2sin^2x

OpenStudy (anonymous):

so is it sin x cos2x - sin3x + cos3x?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

change sin2x into 2sinxcosx and change cos2x into 1-2(sinx)^2